Draw the best Lewis structure for ICl5.
Answer #1

uploaded imageThe VESPR shape of this molecule is octahedral, with one Cl at the bottom and 4 evenly spaced around the I at the middle. I can accept more than an octet in electrons (12), so it has a lone pair up at the top.

Each Cl atom should have 3 pairs of electrons around it, since Cl DOES follow the octet rule. 3 * 2 = 6, + 2 shared electrons that come from the bond with I.
Answer #2
The VESPR shape of this molecule is octahedral, with one Cl at the bottom and 4 evenly spaced around the I at the middle. I can accept more than an octet in electrons (12), so it has a lone pair up at the top.

_________________________________________________________________

The VESPR shape of this molecule is octahedral, with one Cl at the bottom and 4 evenly spaced around the I at the middle. I can accept more than an octet in electrons (12), so it has a lone pair up at the top.

Each Cl atom should have 3 pairs of electrons around it, since Cl DOES follow the octet rule. 3 * 2 = 6, + 2 shared electrons that come from the bond with I.